I asked this question previously but I mistakenly asked incorrectly. I received a response from a very kind gentleman but I need to clarify the question. It goes as follows (my apology);
Brittany plans to take Pre-Calculus, English 12, Physics, Physical Education, Economics, and Anthropology in her senior year. She hopes to have Pre-Calculus and Physics in the morning. What is the probability that her first two classes will be Pre-Calculus or Physics? Stephanie solves the problem incorrectly. In which line does she FIRST make an error?
line 1: P(6,2) = 6!/4! = 30
line 2: P(6,4) = 6!/2! =360
line 3: P (6,2) * P(6,4) = 30*360=10,800
line 4: P(6,6) = 6!/0! - 720
line 5: s/s+f = 720/10,800 = 66.7%
A)Line 1; Correction: P(2,2) = 2!/0! = 2
B)Line 2; Correction: P(4,4) = 4!/0! = 24
C)Line 4; Correction: P(6,6) = 6!*6! = 518,400
D)Line 5; Correction: s/s+f = 10,800/720 = 15%

Found 2 solutions by richard1234, ikleyn:
Answer by richard1234(7193)

Not very clear since the "lines" do not have any explanation, but Line 1 appears to be incorrect, since when counting the total number of schedules, the total number of schedules is simply 6! = 720.

You may have typo-ed since 720/10800 is not 66.7%. The solution in general is very hard to understand.

Answer by ikleyn(52781)

.
I asked this question previously but I mistakenly asked incorrectly. I received a response from a very kind gentleman but I need to clarify the question. It goes as follows (my apology);
Brittany plans to take Pre-Calculus, English 12, Physics, Physical Education, Economics, and Anthropology in her senior year. She hopes to have Pre-Calculus and Physics in the morning. What is the probability that her first two classes will be Pre-Calculus or Physics? Stephanie solves the problem incorrectly. In which line does she FIRST make an error?
line 1: P(6,2) = 6!/4! = 30
line 2: P(6,4) = 6!/2! =360
line 3: P (6,2) * P(6,4) = 30*360=10,800
line 4: P(6,6) = 6!/0! - 720
line 5: s/s+f = 720/10,800 = 66.7%
A)Line 1; Correction: P(2,2) = 2!/0! = 2
B)Line 2; Correction: P(4,4) = 4!/0! = 24
C)Line 4; Correction: P(6,6) = 6!*6! = 518,400
D)Line 5; Correction: s/s+f = 10,800/720 = 15%
~~~~~~~~~~~~~~~~~~~~~~~

Line 1 is irrelevant to the problem.
Same with lines 2 and 3.
Line 4 is written incorrectly.
Line 5 is an error.

Line A) is irrelevant.
Line B) is irrelevant.
Line C) is the error.
Line D) is the error.

The problem was formulated incorrectly from the very beginning.

The "solution" text is nonsense.

The post in whole doesn't make sense.


NO ONE CORRECT AND RELEVANT LINE in the whole text.